OpenStudy (anonymous):

complete the pattern 224,112,56,28,__,__

jimthompson5910 (jim_thompson5910):

You're taking half of each number each time to get the next number (eg: 224/2 = 112 and 112/2 = 56) So take half of 28 to get 28/2 = 14 and take half of 14 to get 14/2 = 7 So the next two numbers are 14 and 7